### v1

# Saber: An Efficient Sampling with Adaptive Acceleration and Backtracking Enhanced Remasking for Diffusion Language Model

Abstract
Diffusion language models (DLMs) are emerging as a powerful and promising alternative to the dominant autoregressive paradigm, offering inherent advantages in parallel generation and bidirectional context modeling. However, the performance of DLMs on code generation tasks, which have stronger structural constraints, is significantly hampered by the critical trade-off between inference speed and output quality. We observed that accelerating the code generation process by reducing the number of sampling steps usually leads to a catastrophic collapse in performance. In this paper, we introduce efficient Sampling with Adaptive acceleration and Backtracking Enhanced Remasking (i.e., Saber), a novel training-free sampling algorithm for DLMs to achieve better inference speed and output quality in code generation. Specifically, Saber is motivated by two key insights in the DLM generation process: 1) it can be adaptively accelerated as more of the code context is established; 2) it requires a backtracking mechanism to reverse the generated tokens. Extensive experiments on multiple mainstream code generation benchmarks show that Saber boosts Pass@1 accuracy by an average improvement of 1.9% over mainstream DLM sampling methods, meanwhile achieving an average 251.4% inference speedup. By leveraging the inherent advantages of DLMs, our work significantly narrows the performance gap with autoregressive models in code generation. [^1] Our code is available at https://github.com/zhaoyMa/Saber.
1 Introduction
Diffusion language models (DLMs) have emerged as a promising non-autoregressive alternative in natural language processing (NLP) fields, with inherent advantages in parallel decoding and bidirectional context modeling through iterative denoising processes (Austin et al., 2021a, Ou et al., 2025, Nie et al., 2025, Ye et al., 2025b). Unlike existing autoregressive models (ARMs) that generate text left-to-right (Radford & Narasimhan, 2018, Radford et al., 2019, Brown et al., 2020, Touvron et al., 2023), DLMs can simultaneously refine multiple token positions by progressively unmasking the generation sequence, enabling global planning and iterative refinement (Ye et al., 2025a, Gong et al., 2025). This paradigm is especially compelling for the structured generation tasks like code generation.
Figure 1: Left: Illustration of DLM Sampling. Right: The trade-off of DLM Sampling between inference speed and output quality on HumanEval benchmark.
Despite these potential advantages, DLMs still lag behind ARMs in practical performance, especially for code generation tasks. The fundamental bottleneck lies in the crucial speed-quality trade-off. As shown in Figure 1, in code generation tasks, the mainstream DLM sampling strategy can lead to a sharp drop in Pass@1 accuracy (even exceeding 60%), once it increases parallelism to reduce the sampling steps, making the DLMs nearly unusable. This severe trade-off prevents DLMs from realizing their inherent parallel generation advantages in practice, as the computational savings from fewer steps are offset by a significant drop in quality.
We argue that the root cause of this severe trade-off stems from two fundamental challenges inherent to the standard DLM sampling process: 1) This process exhibits non-uniform difficulty. The complexity of correctly predicting a token varies significantly across the task, generation context, and token position. Therefore, static acceleration strategies per step (such as using a fixed token number or confidence threshold) are suboptimal. They are often overly conservative in simple stages, sacrificing speed, while being overly aggressive in complex stages, significantly degrading quality. 2) This process can easily be susceptible to error propagation. Unlike ARMs, which only decide what the next token is, DLMs must decide both where and what token to generate. An incorrect choice made early in the process, when the contextual information is sparse, becomes permanently ”locked in” and cannot be revised. This initial error corrupts the context of all subsequent steps, leading to a cascade of failures from which the model cannot recover.
In this paper, we propose Saber, namely efficient Sampling with Adaptive acceleration and Backtracking Enhanced Remasking, a novel training-free sampling algorithm designed to address these two fundamental challenges. Specifically, Saber is built on two key strategies: 1) To address non-uniform difficulty, Saber dynamically adjusts the number of tokens generated in parallel at each step, proceeding cautiously in early, context-poor stages and accelerating as more context is established. 2) To counter error accumulation, Saber introduces a lightweight backtracking mechanism. It allows the model to reverse tokens that are identified as likely errors based on newly available context, enabling a self-correction process that improves final output quality. By introducing these two strategies, Saber achieves substantial speedups while enhancing generation quality.
To evaluate the superiority and generalizability of Saber, we conduct extensive experiments on multiple mainstream code generation benchmarks. We have the findings from the following main aspects: 1) Saber achieves the state-of-the-art performance for DLM sampling in code generation, boosting Pass@1 accuracy by an average improvement of 1.9% over mainstream DLM sampling methods while achieving an average inference speedup of 251.4%. 2) We demonstrate that Saber is a model-agnostic training-free sampling method, which shows effectiveness on various DLMs with consistent performance gains. 3) Through a comprehensive ablation study and variants experiments, we validate that both adaptive acceleration and backtracking-enhanced remasking are integral to Saber’s success. As a result, Saber effectively mitigates the speed-quality trade-off, significantly narrowing the performance gap between DLMs and ARMs for code generation.
Figure 2: Motivation Example. Left: (a) Average confidence per step of DLM sampling. RightL (b) An example of the confidence drop for incorrect tokens of DLM sampling, where gray token means the token generated in the future step.
2 Motivation
Saber is motivated by two key insights from detailed analyses of the DLM sampling process, as shown in Figure 2.
Insight 1: Difficulty Decreases over DLM Generation Process.
The task of generating a masked token is not uniformly difficult throughout the DLM Generation process. In the initial steps, the context is sparse, consisting mostly of ‘[MASK]’ tokens and the initial prompt. In this low-information setting, DLMs are highly uncertain and challenging to generate tokens. However, as more tokens are generated in subsequent steps, the contextual information available to DLMs increases substantially. This richer context progressively reduces DLMs’ uncertainty and simplifies the generation of remaining tokens.
As shown in Figure 2(a), the DLMs’ average prediction confidence steadily increases as more of the sequence is generated. This observation strongly motivates the need for an adaptive acceleration strategy. An idea DLM sampler should be cautious when the context is limited and become progressively more aggressive as DLMs’ confidence grows. This allows for a more principled approach to acceleration that maximizes speed without prematurely committing to low-confidence tokens.
Insight 2: Dynamic Context of DLM Generated Tokens.
A significant difference between DLMs and ARMs is the context of generated tokens. In ARMs, the prefix context for any generated token is fixed. However, in DLMs, the context of generated tokens evolves as ‘[MASK]’ tokens are filled in. Therefore, the DLM’s predicted confidence of generated tokens can dramatically change as new information becomes available. For example, a token might be predicted with high confidence based on sparse local context, only to be revealed as a likely error once a more complete global context is established, as depicted in Figure 2(b).
However, traditional DLM sampling methods are irreversible, i.e., once a token is unmasked, the decision is final and cannot be reversed. This makes them highly susceptible to error propagation, where an overconfident early error corrupts the context for all subsequent steps, leading to a cascade of failures. This issue is a primary driver of the ”catastrophic collapse” when attempting parallel decoding, which highlights the necessity of a backtracking remasking mechanism. By allowing DLMs to revise their own predictions, we can mitigate the risk of early error propagation and enable more robust and aggressive parallel generation.
Summary.
These two insights reveal a fundamental limitation of current samplers: their static and irreversible design fails to account for the dynamic nature of both generation difficulty and contextual certainty during the DLM sampling process. Therefore, in this paper, we argue that an effective DLM sampler must address these limitations by both adapting its generation speed to the evolving context and being able to revise its own past decisions to mitigate error propagation.
3 Related Work
In this section, we outline the two most relevant directions and associated papers of this work.
3.1 Diffusion Language Models for Code
The current landscape of language models is dominated by the autoregressive paradigm (Radford & Narasimhan, 2018, Brown et al., 2020, Touvron et al., 2023, Dubey et al., 2024, Guo et al., 2025). However, their strict left-to-right and token-by-token generation process creates a major bottleneck for inference efficiency and inherently limits parallelism (Li et al., 2025). Therefore, a growing body of research for DLMs has emerged (Li et al., 2022a, Austin et al., 2021a, He et al., 2022), which operate through parallel generation and bidirectional context modeling to address the aforementioned constraints. Recently, large-scale DLMs, such as Dream (Ye et al., 2025b), DiffuLLaMA (Gong et al., 2024), and LLaDA (Nie et al., 2025) have demonstrated performance comparable to similar-scale ARMs, making them a highly promising alternative.
The inherent capabilities of DLMs in global planning and iterative optimization make them naturally suited for code generation (Gong et al., 2025, Li et al., 2025). Therefore, the application of DLMs to this domain has become a significant focus of research (DeepMind, 2025, Gong et al., 2025, Xie et al., 2025, Khanna et al., 2025). However, these works mainly focus on the training process of DLM, while Saber is a training-free DLM sampling method and is orthogonal to them.
3.2 Efficient DLM Sampling Methods
The efficiency of DLMs stems from their ability to generate multiple tokens in parallel (Luxembourg et al., 2025, Yu et al., 2025, Hong et al., 2025, Huang et al., 2025). Some work accelerates this process by setting a fixed threshold, such as Fast-dLLM (Wu et al., 2025), WINO (Hong et al., 2025), and EB-Sampler (Ben-Hamu et al., 2025). However, attempting to unmask multiple tokens in each step degrades the final output quality (Li et al., 2025, Zhang et al., 2025, Wu et al., 2025). Moreover, ReMDM (Wang et al., 2025) proposes a phased sampler that can remask the generated tokens during one of the generation phases. However, the aforementioned methods do not perform well on code generation tasks.
To the best of our knowledge, we are the first to combine adaptive acceleration and backtracking enhanced remasking to achieve improvement for both inference speed and output quality in DLM sampling.
4 Saber
[…]
4.1 Preliminaries
Let a token sequence of length $L$ be denoted by $x=(x_{1},\dots,x_{L})$, where each token $x_{i}$ belongs to a vocabulary $\mathcal{V}$. In the diffusion process, we use a special token [MASK]. At any denoising step $t$, the sequence $x_{t}$ consists of a set of unmasked tokens at indices $\mathcal{U}_{t}$ and a set of masked tokens at indices $\mathcal{M}_{t}$. The DLM $p_{\theta}$, parameterized by $\theta$, takes the partially masked sequence $x_{t}$ as input and outputs a probability distribution over the vocabulary for each masked position $i\in\mathcal{M}_{t}$. We define the model’s confidence in its top prediction for a masked token $i$ as $c_{i}$:
$c_{i}:=\max_{v\in\mathcal{V}}p_{\theta}(x_{i}=v\mid x_{t}),$ (1)
where the mainstream DLM sampling method is to greedily unmask the single token with the highest confidence at each step. Saber improves upon this by the following two key components.
Figure 3: An Overview of Saber in DLM sampling, which consists of two key components, i.e., Adaptive Acceleration via Dynamic Unmasking (AADU) and Backtracking-Enhanced Remasking Mechanism (BERM), during each iterative sampling process.
4.2 Adaptive Acceleration via Dynamic Unmasking
The first component of Saber aims to accelerate inference by unmasking multiple tokens in parallel. Motivated by our observation that the model’s prediction difficulty is non-uniform, we introduce a dynamic and adaptive threshold $\tau_{t}$ to determine which tokens to unmask. This threshold is calculated as the average confidence of all previously unmasked tokens:
$\tau_{t}=\begin{cases}\frac{1}{|\mathcal{U}_{t-1}|}\sum_{j\in\mathcal{U}_{t-1}}c_{j}^{\text{unmask}}&\text{if }t>0,\\ c_{max}&\text{otherwise,}\end{cases}$ (2)
where $c_{j}^{\text{unmask}}$ is the confidence score of token $j$ at the step it was unmasked, and we initialize $\tau_{0}$ to $c_{max}$ for the initial step.
This dynamic threshold naturally encourages a cautious-to-aggressive decoding trajectory. In early steps, when the context is sparse and average confidence is low, $\tau_{t}$ is low, allowing only the most certain tokens to be unmasked. As more high-confidence tokens are generated, $\tau_{t}$ rises, permitting more aggressive parallel unmasking in later, more context-rich stages. Using the threshold $\tau_{t}$, we identify a set of candidate tokens to be drafted, $\mathcal{D}_{t}$, which includes all masked tokens whose current confidence exceeds $\tau_{t}$:
$\mathcal{D}_{t}=\{i\in\mathcal{M}_{t-1}\mid c_{i}\geq\tau_{t}\},$ (3)
where the tokens are provisionally unmasked with their most likely prediction.
Algorithm 1 Pseudocode of Saber in each step.
1:Input: Sequence $x_{t-1}$, DLM $p_{\theta}$, unmasked indices $\mathcal{U}_{t-1}$, unmasked confidences $c^{\text{unmask}}$
2:Output: Updated sequence $x_{t}$
3:// S1: Adaptive Acceleration
4:Compute confidences $c_{i}$ for all $i\in\mathcal{M}_{t-1}$ using $p_{\theta}(\cdot\mid x_{t-1})$.
5:if $t>0$ then
6:  $\tau_{t}\leftarrow\frac{1}{|\mathcal{U}_{t-1}|}\sum_{j\in\mathcal{U}_{t-1}}c_{j}^{\text{unmask}}$.
7:else
8:  $\tau_{t}\leftarrow c_{max}$.
9:end if
[…]
11:Create candidate sequence $x_{t}^{\prime}$ by unmasking tokens in $\mathcal{D}_{t}$.
12:// S2: Backtracking-Enhanced Remasking
[…]
14:Re-evaluate confidences $c_{j}^{t}$ for all $j\in\mathcal{U}_{t-1}$ using the new context $p_{\theta}(\cdot\mid x_{t}^{\prime})$.
15:Initialize an empty set for confidence drops $\Delta$.
16:for each token $j\in\mathcal{U}_{t-1}$ do
17:  $\Delta_{j}\leftarrow c_{j}^{t-1}-c_{j}^{t}$. $\triangleright$ Calculate drop
18:  Add $(j,\Delta_{j})$ to $\Delta$.
19:end for
[…]
21:Create final sequence $x_{t}$ by re-masking tokens at indices $\mathcal{R}_{t}$ in $x_{t}^{\prime}$.
22:Update $c^{\text{unmask}}$ by removing confidences for $j\in\mathcal{R}_{t}$ and adding confidences for $i\in\mathcal{D}_{t}$.
[…]
24:return $x_{t}$, $\mathcal{U}_{t}$, $c^{\text{unmask}}$.
4.3 Backtracking-Enhanced Remasking Mechanism
The second component of Saber introduces a lightweight backtracking mechanism to correct for potential errors made during the aggressive generation in the previous stage. This step is crucial for preventing the error propagation that causes performance collapse.
Unlike methods that may use a fixed threshold, Saber’s backtracking mechanism first determines the number of tokens to revise, $\mu_{t}$, based on how aggressively it generated tokens in the current step:
$\mu_{t}=\max(1,\lfloor|{\mathcal{D}}_{t}|/\mu\rfloor),$ (4)
where $|{\mathcal{D}}_{t}|$ is the size of the newly unmasked set and $\mu$ is a hyperparameter. This ensures that we revise at least one token while limiting the revision to a small fraction of the current step’s output to maintain speed.
Then, we identify which tokens to revise by focusing on those previously unmasked tokens that are most inconsistent with the newly available context. For each existing token $j\in{\mathcal{U}}_{t-1}$, we compute its confidence drop, $\Delta_{j}$, defined as the difference between the unmasked confidences of (t-1)-th time $c_{j}^{t-1}$, and its re-evaluated confidence at the current step $c_{j}^{t}$:
$\Delta_{j}=c_{j}^{t-1}-c_{j}^{t},$ (5)
where a large $\Delta_{j}$ indicates that the model’s belief in its earlier prediction has significantly weakened. We then identify the set of tokens to be reversed, ${\mathcal{R}}_{t}\subseteq{\mathcal{U}}_{t-1}$, by selecting the $\mu_{t}$ tokens that exhibit the largest confidence drop. These are the tokens the model has the most ”regret” about, and they are reverted to [MASK] to be reconsidered in future steps with a richer context.
4.4 Overall
At the conclusion of each step $t$, the final set of unmasked tokens is updated by integrating the outcomes of both the adaptive acceleration and backtracking stages:
$\mathcal{U}_{t}=(\mathcal{U}_{t-1}\cup\mathcal{D}_{t})\setminus\mathcal{R}_{t}.$ (6)
By combining adaptive acceleration with an efficient backtracking mechanism, Saber can decode aggressively while pruning the most probable errors, thus achieving a superior balance between inference speed and generation quality. The pseudocode of Saber in each DLM sampling step is summarized in Algorithm 1.
5 Experiment Setup
In this section, we will provide the setups of our experiments below. The detailed description of experiment setups can be found in Appendix B.
5.1 Datasets
We conduct extensive experiments on five mainstream code generation datasets to demonstrate the effectiveness of Saber, including HumanEval (Chen et al., 2021b): a widely used code generation benchmark consists of 176 Python functions tasks from docstrings, MBPP (Austin et al., 2021b): includs a range of Python programming tasks designed to test basic algorithmic reasoning, HumanEval-ET and MBPP-ET (Dong et al., 2023a): the extended versions of HumanEval and MBPP with 100+ additional test cases, and LiveCodeBench (Jain et al., 2024): a contamination-free benchmark (Dong et al., 2024b) that continuously collects new programming problems from contest platforms (LeetCode, AtCoder, Codeforces).
5.2 Baselines
We conducted a comprehensive evaluation of Saber against existing DLM sampling methods, consisting of: 1) Standard DLM Sampling (Default): In this mode, DLM generates responses by continuously decoding over a predetermined full output length, including confidence-based, entropy-based, and random-based methods. 2) Efficient DLM sampling Methods: Parallelism Increase (p), Semi-autoregressive (SAR) (Nie et al., 2025), WINO (Hong et al., 2025), Fast-dLLM (Wu et al., 2025), and ReMDM (Wang et al., 2025) are recently proposed efficient DLM sampling methods.
5.3 Metric
Our evaluation employs Pass@1 as the primary performance metric, which is calculated as the percentage of problems for which the generated code passes all test cases with a single attempt. The formula is as follows:
$\text{Pass}@1=\frac{1}{|N|}\sum_{i=1}^{|N|}\mathbb{I}(\text{Passed}(\text{Generation}_{i}))$
where $|N|$ is the total number of problems, and the indicator function $\mathbb{I}(\cdot)$ is 1 if the single generation for a given problem passes all its test cases, and 0 otherwise.
In addition to performance, we also measure the Step (i.e., average generation steps per sample) and Time (i.e., total generation time).
5.4 Implementation Details
In this paper, all experiments are conducted on an A6000 GPU (48GB). We employed the LLaDA-8B-Instruct (Nie et al., 2025) as the base model. In the fixed-length scenario, we set the generation length to 256 tokens. For the semi-autoregressive, the block length was configured to 128. All other efficient DLM sampling methods followed the same configuration as their original paper. The default temperature for all baselines is set at 0. To mitigate the instability of the model sampling, we report the average results of five trials in the experiments.
6 Experimental Results
In this section, we present a comprehensive empirical evaluation of Saber. We first compare its performance and efficiency against a wide range of existing DLM sampling methods on multiple code generation benchmarks (§6.1). Next, we demonstrate the model-agnostic nature of Saber by applying it to various state-of-the-art DLMs (§6.2). Finally, we conduct a detailed ablation study to dissect the individual contributions of our proposed components (§6.3) and provide the disscussion of Saber (§6.4).
6.1 Main Results
Table 1 presents the main results of our comparison on the HumanEval, MBPP, HumanEval-ET and MBPP-ET, and LiveCodeBench datasets. The findings clearly demonstrate that Saber sets a new state-of-the-art for DLM sampling in code generation, achieving the highest Pass@1 scores across all benchmarks while simultaneously delivering substantial improvements in inference speed.
Table 1: Comparison of Saber and the existing DLM sampling methods, where the bold indicates the best performance in this column while the underline indicates the second-best performance, and ET means the Pass@1 performance on its extended test case version.
Method HumanEval MBPP LiveCodeBench
Pass@1 $\uparrow$ ET $\uparrow$ Step $\downarrow$ Time $\downarrow$ Pass@1 $\uparrow$ ET $\uparrow$ Step $\downarrow$ Time $\downarrow$ Pass@1 $\uparrow$ Step $\downarrow$ Time $\downarrow$
Standard DLM Sampling
Random 0.1463 0.128 256 1:29:40 0.2295 0.1826 256 2:51:28 0 256 4:09:49
Entropy 0.4146 0.3415 256 1:30:22 0.4215 0.3114 256 2:56:42 0.04 256 4:30:31
Confidence 0.4329 0.3579 256 2:11:52 0.4286 0.3138 256 3:12:08 0.0975 256 5:59:07
Efficient DLM Sampling
Confidence (p=2) 0.3476 0.2866 128 51:13 0.4075 0.2857 128 1:35:13 0.0925 128 2:57:16
SAR (p=2) 0.3598 0.2927 128 1:33:00 0.4005 0.2786 128 1:36:05 0.095 128 2:57:17
Fast-dLLM 0.3963 0.3415 256 59:40 0.4403 0.3044 256 2:30:24 0.0875 256 2:33:29
Fast-dLLM (+parallel) 0.3963 0.3354 96.24 25:25 0.3934 0.2763 73.13 43:18 0.023 96.28 43:22
ReMDM 0.2073 18.29 128 1:26:50 0.3162 0.2248 128 1:28:51 0.033 128 2:50:23
WINO 0.4024 0.3171 100.12 57:10 0.4309 0.3138 88.49 1:44:51 0.0925 77.43 2:40:30
Saber 0.4512 0.3598 118.92 41:55 0.4473 0.3302 110.96 1:33:33 0.11 122.47 2:33:17
Saber Effectively Mitigates the Speed-Quality Trade-off. Compared to standard DLM sampling strategies (Random, Entropy, Confidence), Saber delivers vastly superior performance. For instance, on HumanEval, Saber improves the Pass@1 score from 43.3% (Confidence) to 45.1% while reducing the inference time by nearly 70% (from over 2 hours to just 41 minutes). This result directly refutes the notion that acceleration must come at the cost of quality. While naively increasing parallelism by generating more tokens per step (e.g., Confidence p=2) leads to a significant performance drop (from 43.3% to 34.8%), Saber’s intelligent sampling process successfully avoids this collapse.
Saber Outperforms SOTA Efficient DLM Sampling Methods. When compared to recent efficient sampling methods, Saber establishes a new Pareto frontier for the speed-quality trade-off. WINO, a strong baseline, achieves impressive speed by minimizing decoding steps. However, Saber is even faster in terms of time on most benchmarks, indicating a more efficient computation per step. For example, on HumanEval, Saber is over 25% faster than WINO while also achieving a $\sim$5% higher Pass@1 score. This superior performance is attributed to our lightweight backtracking mechanism, which provides a safety net for the adaptive acceleration, allowing for aggressive parallelization without sacrificing accuracy. Similarly, while Fast-dLLM shows competitive results on MBPP, Saber matches its quality while being nearly 40% faster. On LiveCodeBench, a benchmark designed to be robust against contamination, Saber also achieves the state-of-the-art performance, demonstrating its strong generalization capabilities.
Overall, these results confirm that Saber successfully breaks the existing speed-quality compromise in DLM sampling for code generation.
6.2 Generalizability Across Different DLMs
To validate the “model-agnostic” claim of Saber, we applied it to three distinct open-source DLMs: LLaDA-8B-Instruct (Nie et al., 2025), Dream-v0-Instruct-7B (Ye et al., 2025b), and DiffuCoder-7B-cpGRPO (Gong et al., 2025). We compare the performance of Saber against the standard confidence-based sampler for each DLM on the HumanEval benchmark.
Table 2: Effectiveness of Saber compared to the mainstream DLM sampling method based on different DLMs.
Pass@1 $\uparrow$ Steps $\downarrow$ Time $\downarrow$
LLaDA-8B-Instruct
Confidence (p=1) 0.4329 256 2:11:52
Saber 0.4512 118.92 41:55
Dream-v0-Instruct-7B
Confidence (p=1) 0.2805 256 1:16:15
Saber 0.2927 156.68 46:39
DiffuCoder-7B-cpGRPO
Confidence (p=1) 0.5671 256 1:12:47
Saber 0.5732 140.34 37:08
As shown in Table 2, Saber consistently improves both accuracy and efficiency across all tested models, demonstrating that its benefits are not tied to a specific architecture or training process. For each model, Saber delivers a higher Pass@1 score while simultaneously reducing the number of decoding steps and the total inference time. For instance, on Dream-v0-Instruct-7B, Saber boosts Pass@1 and cuts inference time by nearly 40%. On DiffuCoder-7B, a model specifically optimized for code, Saber further enhances its performance while halving the inference time. This robust performance across different model families validates that Saber addresses fundamental challenges in the DLM sampling process itself, making it a truly general, plug-and-play enhancement.
6.3 Ablation Study
To understand the individual contributions of the two core components of Saber, i.e., Adaptive Acceleration via Dynamic Unmasking and Backtracking-Enhanced Remasking Mechanism, we conducted a thorough ablation study on the HumanEval dataset. The results are presented in Table 3.
Adaptive Acceleration is the Primary Driver of Efficiency. When we remove Adaptive Acceleration via Dynamic Unmasking, the sampler relies only on the backtracking mechanism. While the Pass@1 score remains high at 44.5%, the number of decoding steps reverts to the baseline 256, and the inference time increases dramatically to over 90 minutes. This clearly demonstrates that the adaptive acceleration component is the main source of Saber’s speedup.
Table 3: Ablation study of different components in Saber.
Method Pass@1 $\uparrow$ Steps $\downarrow$ Time $\downarrow$
Ours 0.4512 118.92 41:55
w/o Adaptive Accelerate 0.4451 256 1:32:33
w/o Backtracking Remask 0.3523 65.67 28:30
w/o both 0.3476 128 51:13
$\Delta$ confidence from init. 0.4207 121.46 42:32
Backtracking is Essential for High Quality. Conversely, when we remove Backtracking-Enhanced Remasking Mechanism, the sampler becomes a purely aggressive adaptive accelerator. This variant is extremely fast, finishing in under 30 minutes with only 65.67 steps on average. However, this speed comes at a steep price: the Pass@1 score drops significantly from 45.1% to 35.23%. This result highlights that aggressive parallelization without a corrective mechanism is prone to error propagation, confirming that the backtracking stage is crucial for maintaining high generation quality.
Synergy of Components. Saber achieves the best of both worlds, i.e., a high Pass@1 score of 45.1% and a fast inference time of $\sim$41 minutes, which also shows that the two components are synergistic. The adaptive acceleration allows for aggressive sampling, while the backtracking mechanism provides the necessary safety net to prune errors, enabling a combination of speed and accuracy that neither component can achieve alone. We also validated our dynamic thresholding strategy by replacing it with the average threshold of init generation of tokens ($\Delta$ confidence from init.). This resulted in a lower Pass@1 score of 42.1%, confirming the benefits of an adaptive approach that adjusts to the evolving context.
6.4 Qualitative Analysis
Figure 4 presents a side-by-side comparison of code generated by the default LLaDA sampler and Saber on two problems from the HumanEval benchmark. These examples highlight how Saber’s ability to self-correct prevents the kind of logical failures that plague standard irreversible samplers.
In Problem 1, the default sampler produces code, which is syntactically plausible but logically nonsensical. In contrast, Saber generates the correct, standard nested loop structure. This suggests that the iterative refinement process, guided by backtracking, helps enforce logical and structural coherence, which is paramount in code generation. In Problem 2, the default sampler fundamentally misunderstands the problem’s constraints. Saber, however, correctly decomposes the problem into its core logical components: checking the array’s length and verifying the occurrence count of the maximum element. This ability to correctly construct multi-step, constraint-based logic is a direct benefit of the backtracking mechanism. We hypothesize that the model may initially draft a simpler, incorrect solution, which is then revised in subsequent steps as the evolving context makes the error more apparent, leading to the robust final code.
7 Conclusion
In this paper, we addressed the critical speed-quality trade-off for DLM sampling in code generation and introduced Saber, a novel, training-free sampling algorithm for DLM sampling that combines both adaptive acceleration and backtracking-enhanced remasking mechanism. Our extensive experiments indicate that Saber substantially outperforms existing DLM sampling methods, significantly narrowing the performance gap with autoregressive models in code generation.
Appendix A Case Study
Figure 4: Case Study.
Appendix B Detailed Experimental Setup
B.1 Datasets
We conduct experiments on five code generation datasets to demonstrate the effectiveness of Saber, including HumanEval (Chen et al., 2021b), MBPP (Austin et al., 2021b), HumanEval-ET and MBPP-ET (Dong et al., 2023a), and LiveCodeBench (Jain et al., 2024). For all datasets, tasks are presented in a zero-shot format.
  • HumanEval is a widely used benchmark for evaluating LLMs’ ability to generate correct Python functions from docstrings.
  • MBPP (Mostly Basic Python Problems) consists of small-to-medium Python programming tasks designed to test basic algorithmic reasoning.
  • LiveCodeBench is a contamination-free benchmark that continuously collects new programming problems from contest platforms (LeetCode, AtCoder, Codeforces) and focuses beyond simple code generation to broader code reasoning capabilities.
  • HumanEval-ET and MBPP-ET are extended versions of the original HumanEval and MBPP. They augment each task with over 100 additional test cases and include edge-case tests, which enhances the reliability of the evaluation.
B.2 Baselines
We conducted a comprehensive evaluation of Saber against established baseline decoding methods for DLMs. The results confirm that Saber achieves superior performance, effectively validating its effectiveness.
  • Fixed-length (Default): In this mode, DLM generates responses by continuously decoding over a predetermined full output length. The decoding methods include confidence-based, entropy-based, and random approaches.
  • Semi-autoregressive (SAR): This strategy decodes in blocks from left to right. It thus combines aspects of autoregressive order with diffusion’s simultaneous updates. Within each block, tokens are decoded based on confidence.
  • Parallelism Increase (p), WINO (Hong et al., 2025), Fast-dLLM (Wu et al., 2025), and ReMDM (Wang et al., 2025) are recently proposed efficient DLM sampling methods.
B.3 Metric
Our evaluation employs pass@1 as the primary metric. It is calculated as the percentage of problems for which the generated code passes all test cases with a single attempt. The formula is as follows:
$\text{pass}@1=\frac{1}{|N|}\sum_{i=1}^{|N|}\mathbb{I}(\text{Passed}(\text{Generation}_{i}))$
where $|N|$ is the total number of problems, and the indicator function $\mathbb{I}(\cdot)$ is 1 if the single generation for a given problem passes all its test cases, and 0 otherwise.
In addition to pass@1, we also measure the average decoding steps per sample and the total generation time of each method.
B.4 Implementation Details
In this paper, all experiments were conducted on a workstation equipped with 8 NVIDIA A6000 GPUs (48GB each) and 1TB RAM. We employed the LLaDA-8B-Instruct (Nie et al., 2025) as the base model. In the fixed-length scenario, we set the generation length to 256 tokens. For the semi-autoregressive, the block length was configured to 128. All other efficient DLM sampling methods followed the same configuration as their original paper. The default temperature for all baselines is set at 0. To mitigate the instability of the model sampling, we report the average results of five trials in the experiments.
Appendix C Limitation
Our work has the following two main limitations. First, Saber demands slightly more computational resources than direct sampling in a DLM sampling step. However, compared to the enormous computational overhead of DLMs and our smaller total number of steps, it is marginal and acceptable. Second, we only explore the choice of hyperparameters within reasonable ranges, considering the trade-off between performance and speed, in the right of Figure 1. There is still room for further adjustment of hyperparameters.
Appendix D More Related Works
D.1 Code Generation
Since the advent of artificial intelligence in the 1950s, code generation has been considered the Holy Grail of computer science research (Gulwani et al., 2017). With the rapid expansion of codebases and the increasing capacity of deep learning models, using deep learning for program generation has shown great potential and practicality (Raychev et al., 2014, Ling et al., 2016, Dong et al., 2024a; 2025a, Jiang et al., 2024b; 2025). In recent years, the rise of pre-training techniques has brought new momentum to the field of code generation. For example, studies like CodeT5 (Wang et al., 2021) and UniXcoder (Guo et al., 2022) pre-train models for code generation tasks. With the continual increase in model parameters, researchers have discovered emergent phenomena in LLMs, leading to new breakthroughs . Against this backdrop, LLMs such as AlphaCode (Li et al., 2022b), Codex (Chen et al., 2021a), Starcoder (Li et al., 2023), CodeLlama (Rozière et al., 2023), and DeepSeek Coder (Guo et al., 2024) have emerged.
D.2 Promising Architecture for Language Modeling
While the Transformer has been the foundational architecture for modern language models (Vaswani et al., 2017), the field is experiencing a significant shift with the rise of new paradigms (Dong et al., 2024c; 2025b). Mamba (Gu & Dao, 2023), leveraging a selective State Space Model, presents a compelling alternative that scales linearly with sequence length, effectively overcoming the quadratic complexity bottleneck of Transformers in long-context scenarios (Gu et al., 2021). Simultaneously, a fundamentally different approach is being explored with Diffusion models, which move away from traditional autoregressive generation (Li et al., 2022a). By learning to denoise a sequence from a random state, these models offer a unique framework for highly controllable and iterative text synthesis, signaling a potential new direction for generative AI.